OpenStudy (anonymous):

What is the derivative of this function and what is the normal slope at x=5

OpenStudy (anonymous):

\[x-\sqrt{(x-1)}\]

OpenStudy (anonymous):

$$f(x)=x-\sqrt{x-1}\\f'(x)=1-\frac1{2\sqrt{x-1}}$$

OpenStudy (anonymous):

I keep trying it but i don't get the answer, I'm using the product rule

OpenStudy (anonymous):

Now recall that the normal slope is that of the line perpendicular to our tangent...

OpenStudy (anonymous):

Just use chain rule:$$\frac{d\sqrt{x-1}}{dx}=\frac{d\sqrt{x-1}}{d(x-1)}\times\frac{d(x-1)}{dx}=\frac1{2\sqrt{x-1}}\times1=\frac1{2\sqrt{x-1}}$$

OpenStudy (anonymous):

i was using both

OpenStudy (anonymous):

@Zarkon please help